# Concept-TRAK: Understanding how diffusion models learn concepts through concept-level attribution

### v1

@@ 5.1 Attribution by Customization (AbC)
Following the setup in Wang et al. [47], we use the Attribution by Customization Benchmark (AbC), which evaluates attribution methods on models fine-tuned using one or a few exemplar images to learn new concepts with a special token $\langle V\rangle$. The AbC benchmark measures whether data attribution for images generated using this special token successfully retrieves the exemplars. This setup offers a rare source of ground truth: generated outputs are known to be directly influenced by the exemplars. While this setting lacks generality for large-scale training regimes, it remains the most reliable way to evaluate concept-level attribution in current text-to-image models.
@@
Given that the AbC Benchmark provides ground truth, we frame the problem as a retrieval task and report Recall@10—the proportion of times the exemplar images are successfully retrieved from a pool of 100K LAION images. We compare our method against TRAK [29], D-TRAK [52] and the unlearning-based attribution method (Unlearn) [48]. While the original benchmark involves not only learning examplar using special tokens but also fine-tuning the model’s parameter on the exemplar dataset, real-world use cases more commonly involve investigating the concepts generated or utilized by a single pretrained model. To better reflect this, we adopt textual inversion [11] with a frozen base model (SD1.4v) [34], which only train with a special token $\langle V\rangle$, without parameter update. Further implementation details are provided in the Appendix D.
@@ 6 Related Work
Data Attribution Established data attribution methods include influence functions [32], which approximate leave-one-out retraining via gradients. TRAK and LoGra [29, 6] improve scalability through random projections. Game-theoretic approaches like Data Shapley [21, 14], based on Shapley values [37], were initially limited by retraining costs, but recent work [45, 46] improves efficiency by removing this requirement. Unlearning-based methods [48] offer alternative trade-offs between efficiency and theoretical rigor.
Data Attribution for Diffusion Models Early diffusion attribution methods adapted influence functions [32, 29], but were biased by timestep-dependent gradient norms. Xie et al. [51] addressed this via a re-normalized formulation. Zheng et al. [52] extended TRAK to diffusion models, exploiting its scalability. Lin et al. [23] later proposed the Diffusion Attribution Score, which quantifies per-sample influence by directly comparing predicted distributions, yielding more precise attributions than loss-based approaches.
@@ C.1 Local Concept Attribution
The key insight for local concept attribution is that we should use $p(c_{\text{test}}|x)$ as the reward instead of $p(c|x)$. One approach is to discover $c_{\text{test}}$ through textual inversion [11], enabling attribution for the specific concept as it appears in the generated image rather than the general concept $c$. Preliminary results for this approach are provided in Appendix F.
@@ C.2 External Differentiable Reward Models
It is worth noting that this loss produces gradients equivalent to those from $\nabla$-DB [25] at initialization. However, the derivation and formulation differ significantly: our approach is based on DPS principles, while Liu et al. [25] is derived from GFlowNet foundations [1, 2], resulting in fundamentally different loss functions despite the similar gradient form at initialization.
@@ C.3 Preference Datasets
For preference datasets, one approach would be to first train an explicit reward model and then use the trained reward model to compute utility gradients as described earlier. However, inspired by Direct Preference Optimization (DPO) [33], we can also work with an implicit reward model directly. Here we present a method for measuring utility gradients through an implicit reward formulation.
Given preference pairs $(x^{+},x^{-})$, we can define a Bradley-Terry reward model: $p(x^{+}\succ x^{-})=\sigma(r(x^{+})-r(x^{-}))$, where $x^{+}\succ x^{-}$ means $x^{+}$ is preferred over $x^{-}$, $\sigma$ is the logistic function, and $r(\cdot)$ is the reward model. Following the DPO framework [33], the reward model implied by preference data has the relationship:
@@
In the context of diffusion models, we need to translate this into terms of the denoising score matching loss. For diffusion models, the log probability can be expressed through the DSM loss as $\log p(x)\approx-\mathbb{E}_{t,\epsilon}[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}]$. Therefore, the diffusion-specific DPO loss becomes [44]:
@@
Here, we set the upweight scale $w$ as $1$, since the normalization of the gradient will make the gradient invariant to the upweight scale (Section 3.3). This utility gradient can then be incorporated into our Reward-DPS-Loss, enabling concept attribution with preference data without explicitly training a reward model.
@@ D.1 AbC Benchmark
To address more realistic data attribution scenarios, we modify the original AbC benchmark setup. Rather than fine-tuning model parameters on customization data, we freeze the base model parameters and train only special tokens through textual inversion [11]. Following Wang et al. [47], we create 20 special tokens corresponding to 20 customization concepts. For each special token, we generate 20 images, resulting in 400 total generated images for data attribution evaluation.
@@
Both TRAK and D-TRAK need to specify a regularization hyperparameter $\lambda$. To be more specific, in TRAK [29], we approximate the inverted projected Hessian as $\textbf{H}_{P}^{-1}\approx(\textbf{F}_{P}+\lambda I)^{-1}$, where $\textbf{F}_{P}=\frac{1}{N}\sum_{k}G^{\top}G$ and $G_{ij}=\nabla_{\theta_{j}}L(x_{i};\theta)$. The regularization $\lambda$ is applied to make sure to $\textbf{H}_{P}\approx\textbf{F}_{P}+\lambda I$ is invertible in practice. On the other hand, this regularization makes TRAK-based data attribution effectively ignore components with small eigenvalues, significantly impacting attribution performance [6].
Previous work recommends $\lambda^{*}=0.1\times\text{mean}(\text{eigenvalues}(\textbf{F}_{P}))$ [15]. For a fair comparison, we perform a hyperparameter sweep for TRAK and D-TRAK across $\lambda\in[\lambda^{*}\times 10^{-4},\lambda^{*}\times 10^{4}]$ and report the best performance achieved.
@@
Our method focuses on measuring contributions to specific concepts $c$. We use the synthetic image generation prompt as $c^{+}$ (e.g., “a tattoo of $\langle V\rangle$ snake”) and remove only the special token to create $c^{-}$ (e.g., “a tattoo of snake”). We used $c^{+}$ as the base prompt $c$. We set the regularization parameter to the recommended value $\lambda^{*}=0.1\times\text{mean}(\text{eigenvalues}(\textbf{F}_{P}))$ [15] and use 1024 $x_{t}$ from random trajectories sampled from DDIM sampler for gradient computation.

### v2

@@ 5.1 Attribution by Customization (AbC)
Following the setup in Wang et al. (2023b), we use the Attribution by Customization Benchmark (AbC), which evaluates attribution methods on models fine-tuned using one or a few exemplar images to learn new concepts with a special token $\langle V\rangle$. The AbC benchmark measures whether data attribution for images generated using this special token successfully retrieves the exemplars. This setup offers a rare source of ground truth: generated outputs are known to be directly influenced by the exemplars. While this setting lacks generality for large-scale training regimes, it remains the most reliable way to evaluate concept-level attribution in current text-to-image models.
@@
Given that the AbC Benchmark provides ground truth, we frame the problem as a retrieval task and report Recall@10—the proportion of times the exemplar images are successfully retrieved from a pool of 100K LAION images. We compare our method against TRAK (Park et al., 2023a), D-TRAK (Zheng et al., 2024) and the unlearning-based attribution method (Unlearn) (Wang et al., 2024b). While the original benchmark involves not only learning examplar using special tokens but also fine-tuning the model’s parameter on the exemplar dataset, real-world use cases more commonly involve investigating the concepts generated or utilized by a single pretrained model. To better reflect this, we adopt textual inversion (Gal et al., 2022) with a frozen base model (SD1.4v) (Rombach et al., 2022), which only train with a special token $\langle V\rangle$, without parameter update. Further implementation details are provided in the Appendix D.
@@ 6 Related Work
Data Attribution Established data attribution methods include influence functions (Pruthi et al., 2020), which approximate leave-one-out retraining via gradients. TRAK and LoGra (Park et al., 2023a; Choe et al., 2024) improve scalability through random projections. Game-theoretic approaches like Data Shapley (Jia et al., 2019b; Ghorbani and Zou, 2019), based on Shapley values (Shapley et al., 1953), were initially limited by retraining costs, but recent work (Wang et al., 2023a, 2024a) improves efficiency by removing this requirement. Unlearning-based methods (Wang et al., 2024b) offer alternative trade-offs between efficiency and theoretical rigor.
Data Attribution for Diffusion Models Early diffusion attribution methods adapted influence functions (Pruthi et al., 2020; Park et al., 2023a), but were biased by timestep-dependent gradient norms. Xie et al. (2024) addressed this via a re-normalized formulation. Zheng et al. (2024) extended TRAK to diffusion models, exploiting its scalability. Lin et al. (2025) later proposed the Diffusion Attribution Score, which quantifies per-sample influence by directly comparing predicted distributions, yielding more precise attributions than loss-based approaches.
@@ C.1 Local Concept Attribution
The key insight for local concept attribution is that we should use $p(c_{\text{test}}|x)$ as the reward instead of $p(c|x)$. One approach is to discover $c_{\text{test}}$ through textual inversion [Gal et al., 2022], enabling attribution for the specific concept as it appears in the generated image rather than the general concept $c$. Preliminary results for this approach are provided in Appendix F.
@@ C.2 External Differentiable Reward Models
It is worth noting that this loss produces gradients equivalent to those from $\nabla$-DB [Liu et al., 2024] at initialization. However, the derivation and formulation differ significantly: our approach is based on DPS principles, while Liu et al. [2024] is derived from GFlowNet foundations [Bengio et al., 2021, 2023], resulting in fundamentally different loss functions despite the similar gradient form at initialization.
@@ C.3 Preference Datasets
For preference datasets, one approach would be to first train an explicit reward model and then use the trained reward model to compute utility gradients as described earlier. However, inspired by Direct Preference Optimization (DPO) [Rafailov et al., 2023], we can also work with an implicit reward model directly. Here we present a method for measuring utility gradients through an implicit reward formulation.
Given preference pairs $(x^{+},x^{-})$, we can define a Bradley-Terry reward model: $p(x^{+}\succ x^{-})=\sigma(r(x^{+})-r(x^{-}))$, where $x^{+}\succ x^{-}$ means $x^{+}$ is preferred over $x^{-}$, $\sigma$ is the logistic function, and $r(\cdot)$ is the reward model. Following the DPO framework [Rafailov et al., 2023], the reward model implied by preference data has the relationship:
@@
In the context of diffusion models, we need to translate this into terms of the denoising score matching loss. For diffusion models, the log probability can be expressed through the DSM loss as $\log p(x)\approx-\mathbb{E}_{t,\epsilon}[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}]$. Therefore, the diffusion-specific DPO loss becomes [Wallace et al., 2024]:
@@
Here, we set the upweight scale $w$ as 111, since the normalization of the gradient will make the gradient invariant to the upweight scale (Section 3.3). This utility gradient can then be incorporated into our Reward-DPS-Loss, enabling concept attribution with preference data without explicitly training a reward model.
@@ D.1 AbC Benchmark
To address more realistic data attribution scenarios, we modify the original AbC benchmark setup. Rather than fine-tuning model parameters on customization data, we freeze the base model parameters and train only special tokens through textual inversion [Gal et al., 2022]. Following Wang et al. [2023b], we create 20 special tokens corresponding to 20 customization concepts. For each special token, we generate 20 images, resulting in 400 total generated images for data attribution evaluation.
@@
Both TRAK and D-TRAK need to specify a regularization hyperparameter $\lambda$. To be more specific, in TRAK [Park et al., 2023a], we approximate the inverted projected Hessian as $\textbf{H}_{P}^{-1}\approx(\textbf{F}_{P}+\lambda I)^{-1}$, where $\textbf{F}_{P}=\frac{1}{N}\sum_{k}G^{\top}G$ and $G_{ij}=\nabla_{\theta_{j}}L(x_{i};\theta)$. The regularization $\lambda$ is applied to make sure to $\textbf{H}_{P}\approx\textbf{F}_{P}+\lambda I$ is invertible in practice. On the other hand, this regularization makes TRAK-based data attribution effectively ignore components with small eigenvalues, significantly impacting attribution performance [Choe et al., 2024].
Previous work recommends $\lambda^{*}=0.1\times\text{mean}(\text{eigenvalues}(\textbf{F}_{P}))$ [Grosse et al., 2023]. For a fair comparison, we perform a hyperparameter sweep for TRAK and D-TRAK across $\lambda\in[\lambda^{*}\times 10^{-4},\lambda^{*}\times 10^{4}]$ and report the best performance achieved.
@@
Our method focuses on measuring contributions to specific concepts $c$. We use the synthetic image generation prompt as $c^{+}$ (e.g., “a tattoo of $\langle V\rangle$ snake”) and remove only the special token to create $c^{-}$ (e.g., “a tattoo of snake”). We used $c^{+}$ as the base prompt $c$. We set the regularization parameter to the recommended value $\lambda^{*}=0.1\times\text{mean}(\text{eigenvalues}(\textbf{F}_{P}))$ [Grosse et al., 2023] and use 1024 $x_{t}$ from random trajectories sampled from DDIM sampler for gradient computation.